# Efficient stripe artefact removal by a variational method: application to light-sheet microscopy, FIB-SEM and remote sensing images

### v1

@@ 3.1 Performance Measures
captures the power of corruptions in comparison to the reference as scalar value in $[0,\infty]$ where larger values correspond to higher quality. MS-SSIM on the other hand is an extension of the popular structural similarity index measure
$\mathrm{SSIM}(u,u^{*})=l(u,u^{*})\cdot c(u,u^{*})\cdot\tilde{s}(u,u^{*})\in[0,1]$ (9)
@@
Lastly, the curtaining metric from [19] measures the presence of stripes. It is based on the same ideas as Fourier filtering methods, see Figure 3 (a), and measures how strongly Fourier coefficients are condensed around a central horizontal line. The range of values is $[0,1]$ with larger values corresponding to better quality, i.e., less stripes. A detailed description can be found in [19]. Since this metric is based on similar ideas as Fourier filter methods, it is biased toward such and comparably sensitive to deviations from ideal stripes. However, in contrast to the prior measures, the curtaining metric does not require a reference image and is tailor-made for characterizing corruptions by stripes.
@@ 3.2 Description of employed stripe removal methods
  • The Fourier based multi-directional stripe remover (MDSR) proposed by Liang et al. [6] was described in Section 2.1. We consider a modification restricting filtering to subimages of direction close to the stripe direction. It uses the NSCT decomposition followed by filtering of the Fourier domain of selected subimages. As mentioned before, the number of directions $n_{\mathrm{dir}}=8$ and depth of decomposition $n_{\mathrm{dec}}=5$ can be fixed for all following results. Processing was performed using the ’Nonsubsampled Contourlet Toolbox’ [38] in Matlab. In particular, the function ’nsctdec’ with default options for generating filter banks was used. Therefore, only the standard deviation $\sigma$ of the Gaussian damping function and the damping fall-off parameter $\sigma_{a}$ remain as free parameters.
  • The variational stationary noise remover (VSNR) proposed by Fehrenbach et al. [35] was briefly mentioned in Section 2.2. This variational approach follows the idea of decomposing the image of stripe artifacts into elementary stripe patterns $\psi_{i}$ convolved with location images $\lambda_{i}$. For our application, we chose the real part of three differently sized Gabor filters, see Figure 8, such that short, medium and long stripes can be detected simultaneously. The corresponding objective function reads
@@
    with free weighting parameters $\alpha_{1},\alpha_{2},\alpha_{3}$, $\epsilon>0$ and
    $\left\lVert q\right\rVert_{1,\epsilon}$ $=\sum_{i,j}\phi_{\epsilon}\left(\sqrt{q_{i,j,1}^{2}+q_{i,j,2}^{2}}\right)$
    $\phi_{\epsilon}(x)$ =$=\begin{cases}\frac{x^{2}}{2\epsilon}&\lvert x\rvert\leq\epsilon,\\ \lvert x\rvert-\frac{\epsilon}{2}&\text{otherwise}.\end{cases}$
    We use the Matlab implementation [39] provided by the author Pierre Weiss. For a sufficient approximation of the optimum, we chose to perform 25000 steps of the PDHGMp optimization algorithm.
@@
  • The variational method (6) has been introduced in Section 2.2 and is an extension of the method proposed by Liu et al. [16]. We will refer to it as "general stripe remover" (GSR) and use the variation
[…]
    for 2D images. Free parameters are $\mu_{1}$ and $\mu_{2}$. An approximation of the optimum is attained using 25000 iteration steps of the PDHGMp.
Our main objective for comparing methods was to achieve outcomes which are visually appealing and optimal in terms of the metrics at the same time. However, the performance measures consider only some aspects of the image quality, e.g., the curtaining metric of a strongly smoothed image would suggest optimal image quality even if the actual image content was destroyed. Therefore, visual inspection is required to confirm numerical results. Initial parameters of the methods were obtained from a coarse grid-search and visual and numerical assessment. Afterwards, parameters were manually fine-tuned towards a reasonably optimal solution, i.e., any significant perturbation in parameters yielded either visual degradation, significant reductions of the quality metric values or both. For MDSR and VSNR the results were not clipped towards values in $[0,1]$ when calculating the metrics to reflect the true outcome of the algorithms. However, for proper visualization clipping was applied. From each 3D image used in the following part, an individual slice was selected with prominent stripe artifacts and structures.
@@ 3.3 Synthetic Data
Synthetic images are obtained by modeling the imaging setup in LSFM and simulating light transport [7] using the python package biobeam [18]. The simulation allows for the retrieval of reference images without stripes by neglecting interactions of light and matter. However, this removes natural light attenuation which is usually visible as an intensity decrease along the illumination direction, see Figures 9 (a-b) and 10 (a-b). Hence, the reference images cannot be interpreted as ideal destriping results and a comparison should be done with caution. For details on the simulation and generation of synthetic images, see Appendix A.
The results shown in Figures 9 and 10 display visually significant improvements in image quality across all stripe removal methods. GSR produces the strongest reduction of stripes including the trailing artifacts behind the spherical body in Figure 9 (e) and inside of individual cells in both examples. VSNR removes most thin stripe artifacts with wide stripes remaining visible in Figure 9 (d) and more artifacts remaining inside of cells. Furthermore, faint smearing artifacts are introduced in cavities in Figure 10 (d). MDSR yields the lowest reduction of stripes in both cases with more visible remainders and stronger smearing artifacts, see Figures 9 (c) and 10 (c). PSNR and MS-SSIM show no significant improvement or deterioration for all results. The degree of corruption is reduced significantly for all methods on the cell cluster image but only for GSR on the dispersed cells.
Figure 9: Synthetic image data and destriping results of a cell cluster. Image (a), reference (b), MDSR, $\sigma=12$, $\sigma_{a}=0.3$ (c), VSNR, $\alpha=(3,5,10)$ (d), GSR, $\mu=(\frac{1}{3},\frac{1}{300})$ (e).
Figure 10: Synthetic image data and destriping results of dispersed cells. Image (a), reference (b), MDSR, $\sigma=24$, $\sigma_{a}=0.3$ (c), VSNR, $\alpha=(3,5,10)$ (d), GSR, $\mu=(\frac{1}{2},\frac{1}{60})$ (e).
Metric Input Reference MDSR VSNR GSR(2D)
@@
Table 1: Quality metrics of the images in Figures 9 and 10. Larger is better. The best result is highlighted in green.
@@ 3.4 LSFM images of biological samples
As real image data we consider two real LSFM images with different appearances. The image used in Figure 11 shows an organoid of cultivated mouse intenstine cells where indiviudal cells are clearly visible and separated. On the other hand, Figure 12 displays a cluster of HeLa cells which has a more homogeneous appearance.
Similar observations to the synthetic image data can be made. GSR displays the strongest reduction in stripes including wide and slightly oblique artifacts. However, for the mouse intestine cuboid some faint smearing artifacts were introduced in structural cavities, see Figure 11 (d), and few wide stripes remain visible in Figure 12 (d). VSNR removes all visible stripes in the HeLa-cluster image but is unable to remove wide stripes for the mouse intestine cuboid, see Figures 12 (b) and 11 (b). In the latter, thin but oblique artifacts are reduced but remain visible and very faint smearing artifacts were introduced similar to the observation for GSR. The results by MDSR are comparable to VSNR on the mouse intestine cuboid. Some smearing artifacts are introduced but there are even more remainders of thin oblique stripes than for VSNR. Nonetheless, there are barely any remaining artifacts in the HeLa-cluster image with very similar appearance to the result by GSR.
Figure 11: Mouse intestine cuboid and results. Image (a), MDSR, $\sigma=12$, $\sigma_{a}=0.3$ (b), VSNR, $\alpha=(3,5,10)$ (c), GSR(2D), $\mu=(\frac{1}{3},\frac{1}{300})$ (d). Corresponding curtaining values are 0.58 (a), 0.84 (b), 0.78 (c) and 0.95 (d).
Figure 12: HeLa-cluster and results. Image (a), MDSR, $\sigma=5$, $\sigma_{a}=0.3$ (b), VSNR, $\alpha=(10,5,1)$ (c), GSR(2D), $\mu=(\frac{1}{3},\frac{1}{300})$ (d). Corresponding curtaining values are 0.24 (a), 0.90 (b), 0.99 (c) and 0.96 (d).
@@ 3.5 Application for other imaging methods
We also consider data by other imaging methods where stripe artifacts are commonly encountered. Our examples include a FIB-SEM image of a tin bronze cast alloy and a Terra MODIS satellite image. Both show much more uniformly thin and long stripes compared to the stripes observed in LSFM.
GSR removes stripes entirely from the Terra MODIS image and leaves only few faint remainders of stripes in the FIB-SEM example. Similar results can be observed for VSNR but remainders in the FIB-SEM image are slightly stronger. On the other hand, MDSR introduces visible artifacts in high contrast areas in the Terra MODIS image despite removing all stripes. Additionally, it leaves visible remainders of wide stripes for the FIB-SEM image where multiple stripes occurred close together.
Figure 13: FIB-SEM example and results. Image (a), MDSR, $\sigma=12$, $\sigma_{a}=0.3$ (b), VSNR, $\alpha=(2,10,2)$ (c), GSR(2D), $\mu=(\frac{2}{5},\frac{1}{300})$ (d). Corresponding curtaining values are 0.18 (a), 0.54 (b), 0.95 (c) and 1 (d).
Figure 14: Terra MODIS example and results. Image (a), MDSR, $\sigma=10$, $\sigma_{a}=0.3$ (b), VSNR, $\alpha=(3,5,5)$ (c), GSR(2D), $\mu=(\frac{7}{30},\frac{1}{300})$ (d). Corresponding curtaining values are 1 (a), 1 (b), 1 (c) and 1 (d).
@@ 3.6 Discussion
The results shown in Section 3 reveal that general stripe remover (GSR) outperforms existing methods such as multi-directional stripe remover (MDSR) and Variational Stationary Noise remover (VSNR). It produces consistently good results in terms of stripe removal quality and retention of image structures while offering adjustability through the weighting parameters $\mu_{1}$ and $\mu_{2}$. On the other hand, MDSR and VSNR also show promising results but cannot keep up with GSR on both synthetic and real image data.
@@
The synthetic data which offer stripe-free reference images showed that only GSR is able to remove stripes of larger widths and slightly oblique artifacts such as the trails in Figure 9. This is caused by the choice of functional (6) where stripes are mainly penalized based on directional differences. Hence, even oblique stripe artifacts are affected despite of a slightly larger penalization. On the contrary, VSNR relies heavily on pre-selected patterns such that the removal of oblique artifacts might require an oblique pattern to work well. Similarly, MDSR assumes that stripe information condenses closely into a horizontal coefficient band in the Fourier domain. Deviations in stripe direction will yield a significantly lower reduction of stripe information. With similar arguments we can explain that GSR performs best on the dispersed cells image in Figure 10 where MDSR and VSNR introduce smearing artifacts and yield a lower reduction of stripes.
@@
Including the metrics in Table 1 into our assessment confirms observations for the most part. The high curtaining values produced by MDSR can be attributed to bias as both rely on the same theory and properties of Fourier space. Furthermore, the lack of improvement of MDSR and VSNR on the dispersed cells image is caused by the introduction of smearing artifacts as previously mentioned. Note, that PSNR and MS-SSIM do not improve across all methods. There are multiple reasons that explain this phenomenon. First, the removal of stripe artifacts results in a noticeable decrease in contrast, specifically in areas which were strongly corrupted. Furthermore, the reference image has an overall higher brightness and more details because light attenuation and interactions with matter are neglected during generation. Hence, a full restoration of the original structures is neither anticipated nor possible. Despite of these problems, the fact that PSNR and MS-SSIM have not declined suggests that the structural improvements by removing stripes has compensated any side effect of the removal process.
@@
For the real image data including LSFM and other applications we observe similar results with GSR producing consistently good results. Specifically, the amount of stripe reduction on the mouse intestine cuboid shown in Figure 11 is impressive as almost all artifacts are removed with minimal remainders. This showcases the strengths of GSR as the image is corrupted by a broad range of different stripe artifacts, including various widths, lengths, intensities and deviations in orientations. However, we observe that there are some problems with introducing smearing artifacts in stripe-like cavities of the structure. This can be avoided by adjusting $\mu_{1}$ and $\mu_{2}$ at the cost of lower stripe reduction.
@@
VSNR develops its full potential on the HeLa-cluster image. The result is close to perfect showing no visible artifacts while retaining good contrast and image details. This is achieved since the image reflects an ideal scenario of application where stripe patterns fit perfectly to the stripe artifacts, structures have low contrast and a uniformly colored background. The latter hides smoothing artifacts that may be introduced in the process. The results by GSR on this image are on par with slightly better contrast and detail preservation but few large and smooth stripes remain visible. This is most likely caused by a lower smoothing of artifacts and structures in comparison to VSNR.
MDSR performs generally worse on the real image data. For the mouse intestine cuboid even slightly oblique but thin stripes remain visible. The FIB-SEM example shows noticeably more artifacts than other methods. Finally, on the Terra MODIS example a large number of artifacts has been introduced, e.g., in the upper left landmass. This is surprising as stripe artifacts are periodic and very thin which should be the ideal setting for such a method. Nonetheless, the high contrast in certain areas appears to produce unforeseen problems. On the other side, MDSR achieves a close to perfect result on the HeLa-cluster image. Using the same arguments as before, this is due to the low contrast and uniform background which hides imperfections in the artifact removal. All observations combined show that compared to its variational competitors MDSR is more sensitive to deviations from the model assumption of stripe artifacts being thin and long. Any deviation of length, width or direction yields a lower concentration of stripe information around the horizontal coefficient band. This hampers successful filtering.
For evaluating the quality of stripe removal the curtaining metric proved to be a valuable addition to visual inspection. While it is biased towards Fourier filtering methods, it successfully conveys the corruption of stripe artifacts and improvements made by destriping. Furthermore, it is a good indicator for measuring stripe removal that does not require a reference image such as PSNR or (MS)-SSIM. However, an evaluation of quality still depends mostly on visual inspection and cannot be directly deduced from the metric itself. This becomes clear when comparing the value of curtaining between the corrupted images. We argue that the mouse intestine cuboid displays the most severe corruptions by stripes in terms of visual impairment, yet curtaining values suggest otherwise.
The use of synthetically generated LSFM images for evaluation of stripe removal was, to the best of our knowledge, not considered before in research. However, it provides a broad range of different structures and produces stripe-free reference images which allow us to study the behaviour of stripe removal in much more detail. More precisely, we are able to explore the full capabilities of stripe removal methods as we can produce "extreme" cases and determine limitations. For example, the dispersed cells image in Figure 10 is a challenging setting as stripes appear only as short segments inside of the small cells. It demonstrates limitations such as inconsistent shape reconstruction, insufficient stripe removal or the introduction of smearing artifacts. Furthermore, synthetic image data allows for the use of quality metrics such as PSNR and MS-SSIM which provide more information to asses the stripe removal performance with higher confidence.

### v2

@@ 3.1 Performance Measures
captures the power of corruptions in comparison to the reference as scalar value in $[0,\infty]$ where larger values correspond to higher quality.
MS-SSIM is an extension of the popular structural similarity index measure
$\mathrm{SSIM}(u,u^{*})=l(u,u^{*})\cdot c(u,u^{*})\cdot\tilde{s}(u,u^{*})$ (9)
@@
Lastly, the curtaining metric from [19] measures the presence of stripes. It is based on the same ideas as Fourier filtering methods, see Fig. 2, and measures how strongly Fourier coefficients are condensed around a central horizontal line. The range of values is $[0,1]$ with larger values corresponding to better quality, i.e., less stripes. Since this metric is based on similar ideas as Fourier filter methods, it is biased toward such and comparably sensitive to deviations from ideal stripes. However, in contrast to the prior measures, the curtaining metric does not require a reference image and is tailor-made for characterizing corruptions by stripes.
@@ 3.2 Description of employed stripe removal methods
  • The Fourier based multi-directional stripe remover (MDSR) proposed by Liang et al. [6] as described in Section 2.1. The number of directions $n_{\mathrm{dir}}=8$ and depth of decomposition $n_{\mathrm{dec}}=5$ are fixed for all following results. Processing was performed using the ’Nonsubsampled Contourlet Toolbox’ [45] in Matlab. In particular, the function ’nsctdec’ with default options for generating filter banks was used. The standard deviation $\sigma$ of the Gaussian damping function and the damping fall-off parameter $\sigma_{a}$ remain as free parameters.
  • The variational stationary noise remover (VSNR) proposed by Fehrenbach et al. [36]. For our application, we chose elementary noise patterns $\psi_{i}$ as the real part of three differently sized Gabor filters, see Fig. 7, such that short, medium and long stripes can be detected simultaneously. The corresponding objective function reads
@@
    with free weighting parameters $\alpha_{1},\alpha_{2},\alpha_{3}$, $\epsilon>0$,
    $\left\lVert q\right\rVert_{1,\epsilon}=\sum_{i,j}\phi_{\epsilon}\left(\sqrt{q_{i,j,1}^{2}+q_{i,j,2}^{2}}\right),\quad\phi_{\epsilon}(x)=\begin{cases}\frac{x^{2}}{2\epsilon}&\lvert x\rvert\leq\epsilon,\\ \lvert x\rvert-\frac{\epsilon}{2}&\text{otherwise}.\end{cases}$
    We use the Matlab implementation [46] provided by the author Pierre Weiss and performed 25000 steps of the PDHGMp optimization algorithm.
@@
  • The variational method (6) in its 2D form which we will refer to as ’general stripe remover’ (GSR). Free parameters are $\mu_{1}$ and $\mu_{2}$. An approximation of the optimum is attained using 25000 iteration steps of the PDHGMp.
Our main objective for comparing methods was to achieve outcomes which are visually appealing and optimal in terms of the metrics at the same time. However, the performance measures consider only some aspects of the image quality, e.g., the curtaining metric of a strongly smoothed image would suggest optimal image quality even if the actual image content was destroyed. Therefore, visual inspection is required to confirm numerical results. Initial parameters of the methods were obtained from a coarse grid-search and visual and numerical assessment. Afterwards, parameters were manually fine-tuned towards a reasonably optimal solution where any significant perturbation in parameters yielded either visual degradation, significant reductions of the quality metric values or both.
For MDSR and VSNR the results were not clipped to values in $[0,1]$ when calculating the metrics to reflect the true outcome of the algorithms. However, for proper visualization clipping was applied. From each 3D image used in the following part, a representative slice with prominent stripe artifacts was selected.
@@ 3.3 Synthetic Data
Synthetic images are obtained by modeling the imaging setup in LSFM and simulating light transport [7] using the python package biobeam [18]. The simulation allows for the retrieval of reference images without stripes by neglecting interactions of light and matter. However, this removes natural light attenuation which is usually visible as an intensity decrease along the illumination direction, see Fig. 8 (a-b) and (f-g). Hence, the reference images cannot be interpreted as ideal destriping results and a comparison should be done with caution. For details on the simulation and generation of synthetic images, see Supplement 1. The results shown in Fig. 8 display visually significant improvements in image quality across all stripe removal methods. GSR produces the strongest reduction of stripes including the trailing artifacts below the spherical body, see Fig. 8 (c). VSNR removes most thin stripe artifacts while wide stripes remain visible in Fig. 8 (d). Furthermore, faint smearing artifacts are introduced in cavities in Fig. 8 (i). MDSR yields the lowest reduction of stripes in both cases with more visible remainders and stronger smearing artifacts, see Fig. 8 (c) and (h). PSNR and MS-SSIM (Table 1) do not change significantly for all methods. The curtaining metric increases significantly for all methods on the cell cluster image but only for GSR on the dispersed cells.
Figure 8: Synthetic image data and destriping results of a cell cluster (top) and dispersed cells (bottom). Input image (a), reference (b), MDSR, $\sigma=12$, $\sigma_{a}=0.3$ (c), VSNR, $\alpha=(3,5,10)$ (d), GSR, $\mu=(\frac{1}{3},\frac{1}{300})$ (e), input image (f), reference (g), MDSR, $\sigma=24$, $\sigma_{a}=0.3$ (h), VSNR, $\alpha=(3,5,10)$ (i), GSR, $\mu=(\frac{1}{2},\frac{1}{60})$ (j).
Metric Input Reference MDSR VSNR GSR
@@
Table 1: Quality metrics of the images in Fig. 8. Larger is better. The best result is highlighted in green.
@@ 3.4 LSFM images of biological samples
We consider two LSFM images with different appearances. Fig. 9 (a-d) shows an organoid of cultivated mouse intestine cells where individual cells are clearly visible and separated. Fig. 9 (e-h) displays a cluster of HeLa cells which has a more homogeneous appearance.
As for the synthetic data, GSR yields the strongest reduction in stripes including wide and slightly oblique artifacts. However, for the mouse intestine cells some faint smearing artifacts were introduced in structural cavities, see Fig. 9 (d), and few wide stripes remain visible in Fig. 9 (h). VSNR removes all visible stripes in the HeLa-cluster but is unable to remove wide stripes for the mouse intestine cells, see Fig. 9 (g) and (c). In the latter, thin but oblique artifacts are reduced but remain visible and very faint smearing artifacts were introduced similar to GSR. The results by MDSR are comparable to VSNR on the mouse intestine cells. Some smearing artifacts are introduced but there are even more remainders of thin oblique stripes than for VSNR, compare Fig. 9 (b-c). In the HeLa-cluster barely any artifacts remain, similar to the result of GSR.
Figure 9: Results on LSFM images. Mouse intestine cells (top) and HeLa-cluster (bottom). In put image (a), MDSR, $\sigma=12$, $\sigma_{a}=0.3$ (b), VSNR, $\alpha=(3,5,10)$ (c), GSR, $\mu=(\frac{1}{3},\frac{1}{300})$ (d), input image (e), MDSR, $\sigma=5$, $\sigma_{a}=0.3$ (f), VSNR, $\alpha=(10,5,1)$ (g), GSR, $\mu=(\frac{1}{3},\frac{1}{300})$ (h). Corresponding curtaining values are 0.58 (a), 0.84 (b), 0.78 (c), 0.95 (d), 0.24 (e), 0.90 (f), 0.99 (g) and 0.96 (h).
@@ 3.5 Application for other imaging methods
We also consider data by other imaging methods where stripe artifacts are commonly encountered. Our examples shown in Fig. 10 include a FIB-SEM image of a tin bronze cast alloy and a Terra MODIS satellite image. Both display more uniform thin and long stripes than the LSFM images.
GSR removes stripes entirely from the Terra MODIS image and leaves only few faint remainders of stripes in the FIB-SEM example. Similar results can be observed for VSNR but remainders in the FIB-SEM image are slightly stronger. MDSR introduces visible artifacts in high contrast areas in the Terra MODIS image despite removing all stripes. Additionally, it leaves visible remainders of wide stripes in the FIB-SEM image where multiple stripes occurred close together.
Figure 10: Results on FIB-SEM (top) and Terra MODIS (bottom) examples. Input image (a), MDSR, $\sigma=12$, $\sigma_{a}=0.3$ (b), VSNR, $\alpha=(2,10,2)$ (c), GSR, $\mu=(\frac{2}{5},\frac{1}{300})$ (d), input image (e), MDSR, $\sigma=10$, $\sigma_{a}=0.3$ (f), VSNR, $\alpha=(3,5,5)$ (g), GSR, $\mu=(\frac{7}{30},\frac{1}{300})$ (h). Corresponding curtaining values are 0.18 (a), 0.54 (b), 0.95 (c), 1 (d), 1 (e), 1 (f), 1 (g) and 1 (h).
@@ 3.6 Discussion
The results shown in Section 3 reveal that GSR outperforms existing methods such as MDSR and VSNR. It produces consistently good results in terms of stripe removal quality and retention of image structures while offering adjustability through the weighting parameters $\mu_{1}$ and $\mu_{2}$.
@@
The synthetic data which offer stripe-free reference images showed that only GSR is able to remove stripes of larger widths and slightly oblique artifacts such as the trails in Fig. 8 (a-e). This is explained by the fact that the directional differences in functional (6) also penalize wide and oblique stripe artifacts. On the contrary, VSNR relies heavily on pre-selected patterns such that the removal of oblique artifacts requires oblique patterns to work well. Similarly, MDSR assumes that stripe information condenses closely into a horizontal coefficient band in the Fourier domain. Deviations in stripe direction will yield a significantly lower reduction of stripe information.
@@
The quality metrics in Table 1 confirm visual observations for the most part. The high curtaining values produced by MDSR can be attributed to bias as both rely on the same theory and properties of Fourier space. Furthermore, the lack of improvement of MDSR and VSNR on the dispersed cells image is caused by the introduction of smearing artifacts. Note that PSNR and MS-SSIM do not improve across all methods. There are multiple reasons that explain this phenomenon. First, the removal of stripe artifacts results in a noticeable decrease in contrast, specifically in areas which were strongly corrupted. Furthermore, the reference image has an overall higher brightness and more details because light attenuation and interactions with matter are neglected during generation. Hence, a full restoration of the original structures is neither anticipated nor possible. Despite of these problems, the fact that PSNR and MS-SSIM have not decreased suggests that structural improvements by stripe reduction have compensated any side effect of the removal process.
@@
For the real image data we observe similar results with GSR producing consistently good results. Specifically, the amount of stripe reduction on the mouse intestine cells shown in Fig. 9 is impressive as almost all artifacts are removed. This showcases the strengths of GSR as the image is corrupted by a broad range of different stripe artifacts, including various widths, lengths, intensities and deviations in orientations. However, we observe that there are some problems with introducing smearing artifacts in stripe-like cavities of the structure. This can be avoided by adjusting $\mu_{1}$ and $\mu_{2}$ at the cost of lower stripe reduction.
@@
VSNR develops its full potential on the HeLa-cluster. The result is close to perfect showing no visible artifacts while retaining good contrast and image details. This is achieved since the image reflects an ideal scenario where stripe patterns perfectly fit assumptions on the stripe artifacts, structures have low contrast and a uniformly colored background. The latter hides smoothing artifacts that may be introduced in the process. The results by GSR on this image are on par with slightly better contrast and detail preservation but few large and smooth stripes remain visible. This is most likely caused by a lower smoothing in comparison to VSNR.
The performance of MDSR is somewhat disappointing. MDSR yields the overall lowest stripe reduction while introducing the most severe artifacts across all examples. In particular, the results on the Terra MODIS example is surprising as the thin and periodic stripes should be the ideal setting for this method. However, the high contrast in certain areas seems to produce unforeseen problems. Despite all criticism, the result on the HeLa-cluster shows that performance similar to VSNR and GSR is achievable under the right circumstances. In general, MDSR is more sensitive to stripes deviating from being thin, long and vertically oriented than its variational competitors which is caused by the corresponding influence on stripe information in the Fourier domain.